OpenStudy (anonymous):

What does (fog)(x) mean? Does it mean that f goes into g For example f(x)=(square root)x-3 and g(x)=2x-5

OpenStudy (unklerhaukus):

\[(f\circ g)(x)=f(g(x))\]

hartnn (hartnn):

f o g (x) means to put g(x) in place of x in f(x)

OpenStudy (unklerhaukus):

\[f(x)=\sqrt x-3 \] \[f(g(x)=\sqrt{2x-5}-3\]

OpenStudy (waleed_imtiaz):

it means the composition of f onto g..... So first put g(x) and then f(x);..... LIke in ur question........ fog(x)=f(g(x))=f(2x-5)...... now this is like f(x)..... where x=2x-5.... So according to ur question..... f(x)=(square root)x-3 .... here..... put x=2x-5.... u will get the answer......

OpenStudy (anonymous):

Thank you! I get it now :)

OpenStudy (waleed_imtiaz):

my pleasure...........!